OpenStudy (anonymous):

Name the property the equation illustrates. 9 + (-9) = 0

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

aye hold on 1 sec ight

OpenStudy (anonymous):

Alright. :)

OpenStudy (jordanloveangel):

Commutative Property of Addition a + b = b + a ex:) 2 + 3 = 3 + 2 Commutative Property of Multiplication a • b = b • a ex:) 2 • ( 3 ) = 3 • ( 2 ) Associative Property of Addition a + ( b + c ) = ( a + b ) + c ex:) 2 + ( 3 + 4 ) = ( 2 + 3 ) + 4 Associative Property of Multiplication a • ( b • c ) = ( a • b ) • c ex:) 2 • ( 3 • 4 ) = ( 2 • 3 ) • 4 Distributive Property a • ( b + c ) = a • b + a • c ex:) 2 • ( 3 + 4 ) = 2 • 3 + 2 • 4 Additive Identity Property a + 0 = a ex:) 3 + 0 = 3 Multiplicative Identity Property a • 1 = a ex:) 3 • 1 = 3 Zero Property a • 0 = 0 ex:) 5 • 0 = 0 which is it? what property? i gave you an example.

OpenStudy (jordanloveangel):

do u kinda understand it ? @brandonfrabel

OpenStudy (anonymous):

Oh hell. Now I do! When I first looked at your reply all it showed were weird symbols and diamonds but I just looked back at it and it shows everything perfectly omg thank you

OpenStudy (jordanloveangel):

haha lol np u still need help LML

OpenStudy (anonymous):

Okay I just did a problem fine but there's this one that I can't seem to find the answer for. I looked at all of my notes and what you provided but IDK, to me it just doesn't fit any of the properties? It's 9 + (-9) = 0

OpenStudy (jordanloveangel):

oh lol hold on m i try

OpenStudy (anonymous):

Oh wait, I think it's INverse property of addition. Is that correct?

OpenStudy (jordanloveangel):

Additive Inverse Property

OpenStudy (jordanloveangel):

so u r right

OpenStudy (anonymous):

Yay, thank you :) Okay, so, 0 + x = x..Wouldn't that be the same one?

OpenStudy (anonymous):

Oh wait no wouldn't it be the Identity Property of Addition?

OpenStudy (jordanloveangel):

i dont think so